# UQE: A Query Engine for Unstructured Databases

### v1

@@ 1 Introduction
Data analysis (Elgendy and Elragal, 2014) is essential for making well founded decisions and enabling businesses and society to function more effectively. Relational databases (Codd, 1990; Ramakrishnan and Gehrke, 2002) and the Structured Query Language (SQL) (Chamberlin, 1998) have delivered huge successes in structured data management and analysis. Typically, such data is collected and organized in a pre-defined schema (Elmasri and Navathe, 2013), where the data properties and relationships have been pre-specified, and downstream analysis is restricted to this schema.
In most real-world applications, however, data exists in unstructured formats, such as images, documents and audio recordings. Without preprocessing such data into structured forms, traditional SQL engines can only support limited queries. Preprocessing, including document entity retreival (Yu et al., 2023) and form understanding (Xu et al., 2020), also require training on downstream tasks given a predefined taxonomy. This naturally motivates the question we consider in this paper:
@@
In the literature, full-text search engines (Gospodnetic et al., 2010) support scalable regexp-matching search on unstructured data, but this becomes infeasible for more complex semantic reasoning queries. Retrieval-Augmented Generation (RAG) (Thorne et al., 2020; Lewis et al., 2020; Gao et al., 2023c) allows question answering on a subset of related data, but is not directly applicable to generic analytical tasks with aggregation and semantic queries that spans over an entire large database. Recent advances in Large Language Models (LLMs) (Anthropic, 2024; Achiam et al., 2023) unlock the ability to perform flexible question answering, especially with recent long-context models (Reid et al., 2024). However, setting aside the cost per query, data analytics can still be challenging for LLMs without fine-tuning (Li et al., 2023) or few-shot demonstrations (Chen, 2022), even given structured tables.
Recently, a promising line of work has considered marrying LLMs with programming framworks (Surís et al., 2023), where logical or arithmetic operations are offloaded to program interpreters (Gao et al., 2023b). A most relevant example for analytics and table understanding tasks is Cheng et al. (2022), which augments classical SQL semantics with LLMs as user-defined functions (UDF). While promising, the execution of such SQL programs that embed LLM calls still requires sweeping over the entire database, which is too costly for large collections of unstructured content. To overcome this barrier, we leverage the synergy between LLMs and programmatic execution to define an Unstructured Query Language (UQL) that augments SQL for flexible semantic queries, with a focus on improving scalability and efficiency.
@@
Figure 1: Illustration of unstructured data analysis defined in Section 2
@@ 2 Problem
  • Table / database: We define a table ${\mathcal{T}}=\{{\mathcal{T}}_{i}\}_{i=1}^{N}$ as an unordered set of $|{\mathcal{T}}|=N$ rows, where each row ${\mathcal{T}}_{i}=[{\mathcal{T}}_{i,1},{\mathcal{T}}_{i,2},\ldots,{\mathcal{T}}_{i,M}]$ is an arrary of $M$ elements such that $M$ is the total number of columns in the table. Each row can consist of elements ${\mathcal{T}}_{i,\cdot}$ of heterogeneous types (e.g., datetime, float, enum) with different modalities (e.g., text, image), while elements in each column ${\mathcal{T}}_{\cdot,j}$ must be of the same format and modality.
@@ 3 Unstructured query language
First, we need to formally define the query language, UQL, that talks to the unstructured databases. The idea of defining a natural query language for unstructured data is not completely new (e.g., Cheng et al. (2022), even though UQL has richer semantics), nor is the specific syntax or design of UQL the main focus of this paper. However, we need to define the scope of queries that the engine can handle, and breakdown the semantic meaning of each clause.
@@ 3.1 UQL semantics
The predicates in WHERE are organized in disjunctive normal form (DNF) with AND and OR syntax, so a user can arbitrarily express predicates over concrete and virtual comlumns.
GROUP BY partitions the table into groups, where rows within each group share the same attributes over the keys being grouped by. UQL allows partitioning over virtual columns via natural lanaguage:
@@ 3.2 UQL queries
[fontsize=]mysql SELECT agent_name, "reason to cancel" FROM airline_customer_service_log WHERE "the customer asked to cancel the flight" ORDER BY ticket_price LIMIT 100
@@ 4 Unstructured Query Engine
One straightforward way to run UQL is to use an interpreter that executes queries imperitavely. Cheng et al. (2022) implements an engine in this form, which is able to handle tables of relatively small size. By analogy, this is similar to executing an SQL program using a linear database scan. While it is a valid approach, the latency and cost are prohibitive and generally prevent scaling to real world scenarios.
@@ 4.1.1 Unbiased estimation for aggregation queries
  • Cluster the embeddings $\left\{\vec{e}_{i}\right\}_{i=1}^{N}$ into $K$ disjoint groups $\left\{C_{k}\right\}_{k=1}^{K},C_{k}\subseteq\left\{1\ldots N\right\}$, where $k$ can be a predefined constant or automatically selected (Yuan and Yang, 2019). Each group has size $|C_{k}|$ and $\sum_{k=1}^{K}|C_{k}|=N$. We use $c:\left\{1\ldots N\right\}\mapsto\left\{1\ldots K\right\}$ to denote the cluster index of each row.
@@
Extension The above estimator can be used for other aggregation operartions such as SUM and AVERAGE, including GROUP BY, and allowing concrete columns as operands as well. However, some aggregations such as MAX does not admit such an estimator. UQE in this case can only provide estimates with greater effort. We discuss limitations in Section 7.
@@ 4.1.2 Online learning for non-aggregation queries
In practice, we aim to identify as many rows as possible that satisfy the given condition while adhering to budget constraints (e.g., the total number of tokens allowed to expense). This can be formulated as an online learning problem: given the token budget (or approximately, the number of LLM calls over individual rows, denoted as $B$), we seek to balance exploration (to better understand the semantic landscape of all rows) with exploitation (to maximize recall). Drawing inspiration from Bayesian optimization, which employs a surrogate model learned on-the-fly to inform sequential decision-making Shahriari et al. (2015); Garnett (2023); Hutter et al. (2011), we use a cheap proxy $\hat{g}$ as a surrogate for $f(\mathcal{T}_{i},\texttt{cond})$. At each step $t$, we re-train $\hat{g}$ with the observed data and select its maximizer to query in the next step. See Algorithm 2. Here we use random noise $\epsilon_{t,i}$ to allow some degree of exploration that decays with $t$. Compared to the typical max innerproduct search method prevalent in RAG systems, we rely on the online learning to adjust the beliefs, instead of solely relying on predefined embedding similarities.
@@ 4.2 Compilation
Classically, the goal of a compiler is to translate a high level program to low-level machine code, maintaining exact execution results with improved execution speed. Such a lowering process is usually accompanied by optimizations e.g., fusing, selecting the optimal instructions, and kernel optimizations. Our goal is similar: we would like to compile high-level UQL into low-level machine code, with the distinction that the "machine" is an LLM, and the "low-level code" is the orchestration of prompt calls to the LLM. Given that the primary bottleck is the LLM API calls, we attempt to maintain the execution semantics while minimizing the cost of LLM calls, as in Figure 3.
@@ 4.2.4 Instantiation of kernels
The last step of compilation is to generate the machine specific code (e.g., x86 assembly code) from the intermediate representations (IR). For UQE, this is the process of generating the LLM-specific prompts. For example, when GPT is deployed as the "machine", a system prompt like "You are a helpful assistant" will be added to the queries. This step also sets the correct context (e.g., the correct structured/unstructured column to associate to, the description of the databaes) for the LLM. When such information is not available, one can also leverage the LLM to provide a good suggestion.
@@ 5 Related work
While the unstructured data analytics engine is relatively new, there are several related works in the context of unstructured data query and analysis. Approaches like pattern or regexp matching (Gospodnetic et al., 2010) is scalable but not feasible for complex semantic reasoning. RAG (Lewis et al., 2020; Gao et al., 2023c) based appraoches rely on the retrieval quality and is not directly suitable for aggregation queries over entire database. LLMs (Anthropic, 2024; Achiam et al., 2023; Reid et al., 2024) depict the ability of table analytics (Fang et al., 2024) to some extent (Chen, 2022; Li et al., 2023), but are still not reliable for large unstructured database analytics yet.
Our work is closely related to neural symobilic approaches for unstructured data analytics. Early attempts in this line aim to design specialized neural archictures with inductive biases (e.g., attention) to capture a particular form of operation (e.g., filtering a list of objects based on a natural language predicate by their attention scores) (Yin et al., 2015; Neelakantan et al., 2015; Andreas et al., 2015). Those differentiable neural “operators” can then be chained together to model more compositional queries, and trained end-to-end using graident descent. Another direction, in line with our work, is to augment symbolic programs with learnable operators parametierzed by neural networks (Chen et al., 2020). Those programs are often modeled as discrete latent variables, which can be hard to optimize. In constrast, UQE levarages predictions from LLMs as supervision to train an efficient proxy query model in an online fashion. Similiar to UQE, some recent work (Cheng et al., 2022; Surís et al., 2023) also adopts LLMs as fuzzy query operators. However, the generated programs treat LLMs as an UDF in a SQL program, which can be very expensive to execute on large databases. Our UQE implements similar but augmented semantics with the focus on the cost efficiency and scalability.  Liu et al. (2024) optimizes a similar query engine from the system perspective like cache optimization and deduplication, while our work mainly considers algorithmic improvements and is considered as an approximate query engine (Mozafari and Niu, 2015). These system and algorithm optimizations are actually orthogonal and can be beneficial to jointly consider both for future works.
In a distantly related topic, text2SQL (Yu et al., 2018b, a; Guo et al., 2019; Scholak et al., 2021) also leverages models talking to databases, but is mainly for semantic parsing purpose. While it also leverages the advances in LLMs (Yin et al., 2020; Gao et al., 2023a; Pourreza and Rafiei, 2023; Sun et al., 2023), the execution is still on pure SQL and thus is not suitable for unstructured databases. There are also works on leveraging formal query languages to better query LLMs (Saeed et al., 2023; Beurer-Kellner et al., 2023), with the focus on controllability of the LLM itself rather than performing analytics on external unstructured data.
@@ 6 Experiments
Table 1: Conditional aggregation results on benchmark datasets. We report the relative error and the average cost per query. *gpt-4o is 50% cheaper than gpt-4-turbo so we double its budget of tokens.
@@
lc-gpt-4-turbo lc-claude-3-opus UQE-claude-3-haiku
IMDB sentiment_positive 49.02% $\pm$ 21.23% 56.05% $\pm$ 14.69% 5.75% $\pm$ 3.43%
Average cost per query \$0.37 \$0.61 \$0.01
ABCD account_access 69.25% $\pm$ 32.82% 27.28% $\pm$ 17.05% 11.75% $\pm$ 9.78%
single_item_query 78.42% $\pm$ 9.36% 23.39% $\pm$ 16.14% 12.32% $\pm$ 10.53%
Average cost per query \$0.38 \$0.63 \$0.01
AirDialog book 47.58% $\pm$ 15.24% 54.40% $\pm$ 13.37% 4.98% $\pm$ 2.26%
no_flight 47.92% $\pm$ 21.62% 53.88% $\pm$ 15.77% 8.78% $\pm$ 8.12%
no_reservation 50.54% $\pm$ 21.86% 57.49% $\pm$ 13.08% 7.23% $\pm$ 5.40%
Average cost per query \$0.21 \$0.36 \$0.01
lc-gpt-4o lc-gpt-4-turbo UQE-gpt-4o
Clevr obj_count < 4 22.46% $\pm$ 19.35% 48.15% $\pm$ 41.52% 9.55 $\pm$ 8.55%
# spheres > 3 35.72% $\pm$ 14.95% 91.09% $\pm$ 20.08% 15.14% $\pm$ 10.71%
Average cost per query \$0.33* \$0.33 \$0.20
@@
  • lc-LLM denotes the long-context LLMs that can directly take a subset of database and a natural language question as input, and produce the desired analysis. We mainly evaluate against several model families, including GPT-4 (Achiam et al., 2023) and Claude-3 (Anthropic, 2024). Of course, when evaluating the lc-LLM based approaches, we use the natural language instead of UQL as the prompt.
  • RAG-based can be applied to some non-aggregation queries, such as semantic retrieval. For the retrieval part we use max innerproduct search (MIPS) on top of the same embeddings that are used by UQE, for a controlled experiment.
Datasets: We evaluate different approaches on common analytical tasks in three widely used application domains. We use the datasets that were previously created for discriminative tasks, as these datasets contain both the unstructured columns and the structured ones (the labels in the corresponding dataset). We then hide these structured label columns and perform analytical tasks on the unstructured columns, where these hidden structured columns will be used to compute the groundtruth. The text based tasks include IMDB (Maas et al., 2011) movie reviews, customer service dialogs including Action-Based Conversations Dataset (ABCD (Chen et al., 2021)) and AirDialog (Wei et al., 2018), and image based Clevr (Johnson et al., 2017) dataset. Please refer to Appendix C.1 for more information.
@@
lc-gpt-4-turbo lc-claude-3-opus MIPS UQE
IMDB sentiment_positive 0.397 $\pm$ 0.041 0.556 $\pm$ 0.066 0.875 0.978 $\pm$ 0.003
Average cost per query \$0.38 \$0.63 \$$\simeq\$0$0.02
ABCD account_access 0.045 $\pm$ 0.033 0.080 $\pm$ 0.023 0.961 0.940 $\pm$ 0.019
single_item_query 0.023 $\pm$ 0.021 0.082 $\pm$ 0.030 0.266 0.935 $\pm$ 0.006
Average cost per query \$ 0.76 \$1.23 \$$\simeq\$0$0.03
AirDialog book 0.327 $\pm$ 0.0667 0.585 $\pm$ 0.025 0.930 0.979 $\pm$ 0.010
no_flight 0.066 $\pm$ 0.037 0.228 $\pm$ 0.068 0.867 0.928 $\pm$ 0.018
no_reservation 0.156 $\pm$ 0.075 0.297 $\pm$ 0.043 0.965 0.969 $\pm$ 0.004
cancel 0.006 $\pm$ 0.009 0.013 $\pm$ 0.008 0.066 0.741 $\pm$ 0.205
Average cost per query \$0.43 \$ 0.74 \$$\simeq\$0$0.01
Clevr obj_count < 4 0.058 $\pm$ 0.026 0.066 $\pm$ 0.023 0.023 0.850 $\pm$ 0.025
# spheres > 3 0.037 $\pm$ 0.027 0.099 $\pm$ 0.023 0.145 0.633 $\pm$ 0.177
Average cost per query \$0.38 \$0.21 \$$\simeq\$0$0.08
@@
Setup: We use voyage-2 (voyage, 2) to embed the text-based unstructured columns, and Vertex (Vertex, ) for multimodal embeddings. For budget constraint queries, we allow different approachces to access at most 128 rows in the database by default.

### v2

@@ 1 Introduction
Data analysis [13] is essential for making well founded decisions and enabling businesses and society to function more effectively. Relational databases [12, 32] and the Structured Query Language (SQL) [7] have delivered huge successes in structured data management and analysis. Typically, such data is collected and organized in a pre-defined schema [14], where the data properties and relationships have been pre-specified, and downstream analysis is restricted to this schema.
In most real-world applications, however, data exists in unstructured formats, such as images, documents and audio recordings. Without preprocessing such data into structured forms, traditional SQL engines can only support limited queries. Preprocessing, including document entity retreival [45] and form understanding [42], also require training on downstream tasks given a predefined taxonomy. This naturally motivates the question we consider in this paper:
@@
In the literature, full-text search engines [20] support scalable regexp-matching search on unstructured data, but this becomes infeasible for more complex semantic reasoning queries. Retrieval-Augmented Generation (RAG) [39, 24, 18] allows question answering on a subset of related data, but is not directly applicable to generic analytical tasks with aggregation and semantic queries that spans over an entire large database. Recent advances in Large Language Models (LLMs) [4, 2] unlock the ability to perform flexible question answering, especially with recent long-context models [33]. However, setting aside the cost per query, data analytics can still be challenging for LLMs without fine-tuning [25] or few-shot demonstrations [9], even given structured tables.
Recently, a promising line of work considered marrying LLMs with programming frameworks [38], where logical or arithmetic operations are offloaded to program interpreters [17]. A most relevant example for analytics and table understanding tasks is Cheng et al. [11], which augments classical SQL semantics with LLMs as user-defined functions (UDF). While promising, the execution of such SQL programs that embed LLM calls still requires sweeping over the entire database, which is too costly for large collections of unstructured content. To overcome this barrier, we leverage the synergy between LLMs and programmatic execution to define an Unstructured Query Language (UQL) that augments SQL for flexible semantic queries, with a focus on improving scalability and efficiency.
@@ 2 Problem
Figure 1: Illustration of unstructured data analysis defined in Section 2.
@@
  • Table / database: We define a table ${\mathcal{T}}=\{{\mathcal{T}}_{i}\}_{i=1}^{N}$ as an unordered set of $|{\mathcal{T}}|=N$ rows, where each row ${\mathcal{T}}_{i}=[{\mathcal{T}}_{i,1},{\mathcal{T}}_{i,2},\ldots,{\mathcal{T}}_{i,M}]$ is an array of $M$ elements such that $M$ is the total number of columns in the table. Each row can consist of elements ${\mathcal{T}}_{i,\cdot}$ of heterogeneous types (e.g., datetime, float, enum) with different modalities (e.g., text, image), while elements in each column ${\mathcal{T}}_{\cdot,j}$ must be of the same format and modality.
@@ 3 Unstructured query language
First, we need to formally define the query language, UQL, that talks to the unstructured databases. The idea of defining a natural query language for unstructured data is not completely new (e.g., Cheng et al. [11], even though UQL has richer semantics), nor is the specific syntax or design of UQL the main focus of this paper. However, we need to define the scope of queries that the engine can handle, and breakdown the semantic meaning of each clause.
@@ 3.1 UQL semantics
The predicates in WHERE are organized in disjunctive normal form (DNF) with AND and OR syntax, so a user can arbitrarily express predicates over concrete and virtual columns.
GROUP BY partitions the table into groups, where rows within each group share the same attributes over the keys being grouped by. UQL allows partitioning over virtual columns via natural language:
@@ 3.2 UQL queries
[fontsize=]mysql SELECT agent_name, "reason to cancel" FROM airline_customer_service_log WHERE "the customer asked to cancel the flight" LIMIT 100
@@ 4 Unstructured Query Engine
One straightforward way to run UQL is to use an interpreter that executes queries imperatively. Cheng et al. [11] implements an engine in this form, which is able to handle tables of relatively small size. By analogy, this is similar to executing an SQL program using a linear database scan. While it is valid, the latency and cost are prohibitive and generally prevent scaling to real world scenarios.
@@ 4.1.1 Unbiased estimation for aggregation queries
  • Cluster the embeddings $\left\{\vec{e}_{i}\right\}_{i=1}^{N}$ into $K$ disjoint groups $\left\{C_{k}\right\}_{k=1}^{K},C_{k}\subseteq\left\{1\ldots N\right\}$, where $k$ can be a predefined constant or automatically selected [48]. Each group has size $|C_{k}|$ and $\sum_{k=1}^{K}|C_{k}|=N$. We use $c:\left\{1\ldots N\right\}\mapsto\left\{1\ldots K\right\}$ to denote the cluster index of each row.
@@
Extension The above estimator can be used for other aggregation operations such as SUM and AVERAGE, including GROUP BY, and allowing concrete columns as operands as well. However, some aggregations such as MAX does not admit such an estimator. UQE in this case can only provide estimates with greater effort. We discuss limitations in Section 7.
@@ 4.1.2 Online learning for non-aggregation queries
In practice, we aim to identify as many rows as possible that satisfy the given condition while adhering to budget constraints (e.g., the total number of tokens allowed to expense). This can be formulated as an online learning problem: given the token budget (or approximately, the number of LLM calls over individual rows, denoted as $B$), we seek to balance exploration (to better understand the semantic landscape of all rows) with exploitation (to maximize recall). Drawing inspiration from Bayesian optimization, which employs a surrogate model learned on-the-fly to inform sequential decision-making [36, 19, 22], we use a cheap proxy $\hat{g}$ as a surrogate for $f(\mathcal{T}_{i},\texttt{cond})$. At each step $t$, we re-train $\hat{g}$ with the observed data and select its maximizer to query in the next step. See Algorithm 2. Here we use random noise $\epsilon_{t,i}$ to allow some degree of exploration that decays with $t$. Compared to the typical max inner-product search method prevalent in RAG systems, we rely on the online learning to adjust the beliefs, instead of solely relying on predefined embedding similarities.
@@ 4.2 Compilation
Classically, the goal of a compiler is to translate a high level program to low-level machine code, maintaining exact execution results with improved execution speed. Such a lowering process is usually accompanied by optimizations e.g., fusing, selecting the optimal instructions, and kernel optimizations. Our goal is similar: we would like to compile high-level UQL into low-level machine code, with the distinction that the "machine" is an LLM, and the "low-level code" is the orchestration of prompt calls to the LLM. Given that the primary bottleneck is the LLM API calls, we attempt to maintain the execution semantics while minimizing the cost of LLM calls, as in Figure 3.
@@ 4.2.4 Instantiation of kernels
The last step of compilation is to generate the machine specific code (e.g., x86 assembly code) from the intermediate representations (IR). For UQE, this is the process of generating the LLM-specific prompts. For example, when GPT is deployed as the "machine", a system prompt like "You are a helpful assistant" will be added to the queries. This step also sets the correct context (e.g., the correct structured/unstructured column to associate to, the description of the databases) for the LLM. When such information is not available, one can also leverage the LLM to provide a good suggestion.
@@ 5 Related work
While the unstructured data analytics engine is relatively new, there are several related works in the context of unstructured data query and analysis. Approaches like pattern or regexp matching [20] is scalable but not feasible for complex semantic reasoning. RAG [24, 18] based approaches rely on the retrieval quality and is not directly suitable for aggregation queries over entire database. LLMs [4, 2, 33] depict the ability of table analytics [15] to some extent [9, 25], but are still not reliable for large unstructured database analytics yet.
Our work is closely related to neural symboilic approaches for unstructured data analytics. Early attempts in this line aim to design specialized neural architectures with inductive biases (e.g., attention) to capture a particular form of operation (e.g., filtering a list of objects based on a natural language predicate by their attention scores) [43, 29, 3]. Those differentiable neural “operators” can then be chained together to model more compositional queries, and trained end-to-end using gradient descent. Another direction, in line with our work, is to augment symbolic programs with learnable operators parameterized by neural networks [10]. Those programs are often modeled as discrete latent variables, which can be hard to optimize. In contrast, UQE leverages predictions from LLMs as supervision to train an efficient proxy query model in an online fashion. Similar to UQE, some recent work [11, 38] also adopts LLMs as fuzzy query operators. However, the generated programs treat LLMs as an UDF in a SQL program, which can be very expensive to execute on large databases. Our UQE implements similar but augmented semantics with the focus on the cost efficiency and scalability.  Liu et al. [26] optimizes a similar query engine from the system perspective like cache optimization and deduplication, while our work mainly considers algorithmic improvements and is considered as an approximate query engine [28]. These system and algorithm optimizations are actually orthogonal and can be beneficial to jointly consider both for future works.
In a distantly related topic, text2SQL [47, 46, 21, 35] also leverages models talking to databases, but is mainly for semantic parsing purpose. While it also leverages the advances in LLMs [44, 16, 31, 37], the execution is still on pure SQL and thus is not suitable for unstructured databases. There are also works on leveraging formal query languages to better query LLMs [34, 6], with the focus on controllability of the LLM itself rather than performing analytics on external unstructured data.
@@ 6 Experiments
Table 1: Conditional aggregation results on benchmark datasets. We report the relative error and the average cost per query. *gpt-4o is 50% cheaper than gpt-4-turbo so we double its budget of tokens; we use claude-3-haiku as the backend LLM.
@@
lc-gpt-4-turbo lc-claude-3-opus UDF UQE
IMDB sentiment_positive 49.02% $\pm$ 21.23% 56.05% $\pm$ 14.69% 13.67% $\pm$ 6.24% 5.75% $\pm$ 3.43%
Average cost per query \$0.37 \$0.61 \$0.01 \$0.01
ABCD account_access 69.25% $\pm$ 32.82% 27.28% $\pm$ 17.05% 18.99% $\pm$ 9.85% 11.75% $\pm$ 9.78%
single_item_query 78.42% $\pm$ 9.36% 23.39% $\pm$ 16.14% 26.95% $\pm$ 22.16% 12.32% $\pm$ 10.53%
Average cost per query \$0.38 \$0.63 \$0.01 \$0.01
AirDialog book 47.58% $\pm$ 15.24% 54.40% $\pm$ 13.37% 10.15% $\pm$ 7.64% 4.98% $\pm$ 2.26%
no_flight 47.92% $\pm$ 21.62% 53.88% $\pm$ 15.77% 21.08% $\pm$ 16.78% 8.78% $\pm$ 8.12%
no_reservation 50.54% $\pm$ 21.86% 57.49% $\pm$ 13.08% 21.19% $\pm$ 12.10% 7.23% $\pm$ 5.40%
Average cost per query \$0.21 \$0.36 \$0.01 \$0.01
lc-gpt-4o lc-gpt-4-turbo UDF-gpt-4o UQE-gpt-4o
Clevr obj_count < 4 22.46% $\pm$ 19.35% 48.15% $\pm$ 41.52% 31.04% $\pm$ 25.15% 9.55 $\pm$ 8.55%
# spheres > 3 35.72% $\pm$ 14.95% 91.09% $\pm$ 20.08% 19.35% $\pm$ 13.81% 15.14% $\pm$ 10.71%
Average cost per query \$0.33* \$0.33 \$0.20 \$0.20
@@
  • lc-LLM denotes the long-context LLMs that can directly take a subset of database and a natural language question as input, and produce the desired analysis. We mainly evaluate against several model families, including GPT-4 [2] and Claude-3 [4]. Of course, when evaluating the lc-LLM based approaches, we use the natural language instead of UQL as the prompt.
  • RAG-based can be applied to some non-aggregation queries, such as semantic retrieval. For the retrieval part we use max inner-product search (MIPS) on top of the same embeddings that are used by UQE, for a controlled experiment.
  • UDF simply treats the LLM calls as User-defined function of an SQL engine, with the same budget as UQE by default. This approach will not have the advanced sampling / search algorithm as used in UQE, which also serves as an ablation for the effectiveness of our UQE.
Datasets: We evaluate different approaches on common analytical tasks in three widely used application domains. We use the datasets that were previously created for discriminative tasks, as these datasets contain both the unstructured columns and the structured ones (the labels in the corresponding dataset). We then hide these structured label columns and perform analytical tasks on the unstructured columns, where these hidden structured columns will be used to compute the ground-truth. The text based tasks include IMDB [27] movie reviews, customer service dialogs including Action-Based Conversations Dataset (ABCD [8]) and AirDialog [41], and image based Clevr [23] dataset. Please refer to Appendix C.1 for more information.
@@
lc-gpt-4-turbo lc-claude-3-opus UDF MIPS UQE
IMDB sentiment_positive 0.397 $\pm$ 0.041 0.556 $\pm$ 0.066 0.505 $\pm$ 0.030 0.875 0.978 $\pm$ 0.003
Average cost per query \$0.38 \$0.63 \$0.02 \$$\simeq\$0$0.02
ABCD account_access 0.045 $\pm$ 0.033 0.080 $\pm$ 0.023 0.076 $\pm$ 0.017 0.961 0.940 $\pm$ 0.019
single_item_query 0.023 $\pm$ 0.021 0.082 $\pm$ 0.030 0.065 $\pm$ 0.017 0.266 0.935 $\pm$ 0.006
Average cost per query \$ 0.76 \$1.23 \$0.03 \$$\simeq\$0$0.03
AirDialog book 0.327 $\pm$ 0.0667 0.585 $\pm$ 0.025 0.342 $\pm$ 0.031 0.930 0.979 $\pm$ 0.010
no_flight 0.066 $\pm$ 0.037 0.228 $\pm$ 0.068 0.144 $\pm$ 0.034 0.867 0.928 $\pm$ 0.018
no_reservation 0.156 $\pm$ 0.075 0.297 $\pm$ 0.043 0.145 $\pm$ 0.042 0.965 0.969 $\pm$ 0.004
cancel 0.006 $\pm$ 0.009 0.013 $\pm$ 0.008 0.013 $\pm$ 0.009 0.066 0.741 $\pm$ 0.205
Average cost per query \$0.43 \$ 0.74 \$0.01 \$$\simeq\$0$0.01
Clevr obj_count < 4 0.058 $\pm$ 0.026 0.066 $\pm$ 0.023 0.093 $\pm$ 0.031 0.023 0.850 $\pm$ 0.025
# spheres > 3 0.037 $\pm$ 0.027 0.099 $\pm$ 0.023 0.089 $\pm$ 0.017 0.145 0.633 $\pm$ 0.177
Average cost per query \$0.38 \$0.21 \$0.08 \$$\simeq\$0$0.08
@@
Setup: We use voyage-2 [1] to embed the text-based unstructured columns, and Vertex [40] for multimodal embeddings. For budget constraint queries, we allow different approachces to access at most 128 rows in the database by default.